Calculate the following integrals by using the integration

methods.

1

a) ∫ (4π‘₯^34 + 8π‘₯^3 + 15π‘₯)𝑑π‘₯ / (√π‘₯^2 + 4x)

0


b) ∫ 𝑑π‘₯ / sin π‘₯ βˆ™ cos^2 π‘₯


1
Expert's answer
2021-07-19T08:14:47-0400

(a)


"u=x+2, du=dx"

"4\ud835\udc65^4 + 8\ud835\udc65^3 + 15\ud835\udc65=4x^3(x+2)+15x"

"=4u(u-2)^3+15(u-2)"

"=4u^4-24u^3+48u^2-32u+15u-30"


"=4u^2(u^2-4)+16u^2-24u(u^2-4)-96u"

"+48u^2-32u+15u-30"

"=4u^2(u^2-4)-24u(u^2-4)+64(u^2-4)"


"-113u+226"




"\sqrt{x^2+4x}=\sqrt{x^2+4x+4-4}=
\sqrt{u^2-4}"

Table of Integrals


"\int u^2\sqrt{u^2-a^2}du=\dfrac{u}{8}(2u^2-a^2)\sqrt{u^2-a^2}"

"-\dfrac{a^4}{8}\ln|u+\sqrt{u^2-a^2}|+C"


"\int u\sqrt{u^2-a^2}du=\dfrac{1}{3}(u^2-a^2)\sqrt{u^2-a^2}+C"

"-\dfrac{a^4}{8}\ln|u+\sqrt{u^2-a^2}|+C"


"\int \sqrt{u^2-a^2}du=\dfrac{u}{2}\sqrt{u^2-a^2}"

"-\dfrac{a^2}{2}\ln|u+\sqrt{u^2-a^2}|+C"

"\int\dfrac{u}{\sqrt{u^2-a^2}}du=\sqrt{u^2-a^2}+C"


"\int\dfrac{1}{\sqrt{u^2-a^2}}du=\ln|u+\sqrt{u^2-a^2}|+C"


"\int\dfrac{4u^4-24u^3+48u^2-17u-30}{\sqrt{u^2-4}}du"




"=\int4u^2\sqrt{u^2-4}du-\int24u\sqrt{u^2-4}du"

"+\int64\sqrt{u^2-4}du-\int\dfrac{113u}{\sqrt{u^2-4}}du+\int\dfrac{226}{\sqrt{u^2-4}}du"

"=u(u^2-2)\sqrt{u^2-4}-8\ln|u+\sqrt{u^2-4}|-"

"-8(u^2-4)\sqrt{u^2-4}"

"+32u\sqrt{u^2-4}-128\ln|u+\sqrt{u^2-4}|"

"-113\sqrt{u^2-4}+226\ln|u+\sqrt{u^2-4}|+C"

"=\sqrt{u^2-4}(u^3-2u-8u^2+32+32u-113)"

"+90\ln|u+\sqrt{u^2-4}|+C"

"=\sqrt{x^2+4x}((x+2)^3-8(x+2)^2+30(x+2)-81)"

"+90\ln|x+2+\sqrt{x^2+4x}|+C"

"\displaystyle\int_{0}^{1}\dfrac{4𝑥^4 + 8𝑥^3 + 15𝑥}{
\sqrt{x^2+4x}}dx"

"=\sqrt{5}(27-72+90-81)+90\ln(3+\sqrt{5})"

"-90\ln(2)=90\ln(\dfrac{3+\sqrt{5}}{2})-36"


(b)


"\int\dfrac{dx}{\sin x\cos^2 x}=\int\dfrac{(1+\tan^2 x)dx}{\sin x}"

"\int\dfrac{dx}{\sin x}=\int\dfrac{\sin x dx}{\sin^2 x}=\int\dfrac{\sin xdx}{1-\cos^2 x}"

"=\dfrac{1}{2}\int\dfrac{\sin xdx}{1-\cos x}-\dfrac{1}{2}\int\dfrac{\sin xdx}{1+\cos x}"

"=\dfrac{1}{2}\ln(1-\cos x)-\dfrac{1}{2}\ln(1+\cos x)+C_1"


"\int\dfrac{\tan^2 xdx}{\sin x}=\int\dfrac{\sin x dx}{\cos^2 x}=\dfrac{1}{\cos x}+C_2"

Therefore


"\int\dfrac{dx}{\sin x\cos^2 x}=\dfrac{1}{2}\ln(1-\cos x)-\dfrac{1}{2}\ln(1+\cos x)"

"+\dfrac{1}{\cos x}+C"